OpenStudy (amy0799):

if f(x) = (x^2-c^2)/(x^2+c^2) where c is a constant, find f'(x)

OpenStudy (irishboy123):

\[f(x) = \frac {x^2-c^2}{x^2+c^2}\] simplify? \[f(x) = \frac {x^2+c^2-2c^2}{x^2+c^2} = 1 - \frac{2c^2}{x^2 + c^2}\] can you finish this?

OpenStudy (amy0799):

i thought it would be \[\frac{ 2x-2c }{ 2x+2c }\]

OpenStudy (irishboy123):

why did you think that?

OpenStudy (amy0799):

hold on, i know what i did wrong

OpenStudy (irishboy123):

and remember, c is a constant so \(\frac{d}{dx} \left[ c^2 \right] = 0\)

OpenStudy (amy0799):

\[\frac{ (x ^{2} +c ^{2})2x-(x ^{2{}}-c ^{2})2x}{ (x ^{2} +c ^{2})^{2}}\] is this right?

OpenStudy (irishboy123):

you are applying the quotient rule. before i look at your work, is that what you are supposed to be doing with this because it is a silly way to do it. it just adds complications. let me know either way and we can proceed

OpenStudy (irishboy123):

your application of the quotient rule is correct

OpenStudy (irishboy123):

if you wish to do it this way, next step is to simplify the numerator

OpenStudy (amy0799):

simpliflying it would get me \[\frac{ 4xc ^{2} }{ (x ^{2}+c ^{2})^{2} }\]

OpenStudy (irishboy123):

well done!

OpenStudy (amy0799):

so that's the answer?

OpenStudy (irishboy123):

yes

OpenStudy (amy0799):

thank you!

OpenStudy (irishboy123):

mp